40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 21 18:40:46 2014, manasa, Update, General, End laser fiber setup 
    Reply  Fri Oct 24 10:08:24 2014, manasa, Update, General, Y AUX laser - fiber coupled 
       Reply  Fri Oct 24 18:53:23 2014, rana, Update, General, Y AUX laser - fiber coupled 
       Reply  Mon Oct 27 12:17:46 2014, manasa, Update, General, Y AUX laser coupling telescope design AUXmode.zipAUXmode.png
          Reply  Wed Oct 29 11:45:11 2014, manasa, Update, General, Y AUX laser - fiber coupled (52%) 31.png
Message ID: 10629     Entry time: Tue Oct 21 18:40:46 2014     Reply to this: 10638
Author: manasa 
Type: Update 
Category: General 
Subject: End laser fiber setup 

 [Manasa, Diego]

This is our first time touching tables for Frequency Offset Locking. 

The goal was to couple the 1064nm that leaks after the SHG crystal and couple it into the fiber before we run it along the length of the arm. 

The fiber has been mounted at the end but there is no light coupled into the fiber as yet.

In the process, the following were done:

1. ETMY oplev servo disabled. This was enabled after the work.
2. NPRO laser power was reduced so that nothing gets burnt accidently while putting things on the table. This was also reset after the work.

The arms could be locked to green and IR after the work. So I am hoping today's work will not affect locking.

ELOG V3.1.3-